jigglypuff314 (jigglypuff314):

ABC + DEF + GHI = JJJ A to I are all different numbers/digits. What is J? (no leading zeros allowed)

OpenStudy (primeralph):

You know the magic word.

jigglypuff314 (jigglypuff314):

pwease?

OpenStudy (primeralph):

Nope. It's a song actually.

jigglypuff314 (jigglypuff314):

idk :/ :3

OpenStudy (primeralph):

What a shame.

jigglypuff314 (jigglypuff314):

jigglypuff? brb me has dinner

OpenStudy (primeralph):

You're not the real jigglypuff if you don't know the song.

jigglypuff314 (jigglypuff314):

hi I'm back ^.^ and I know I'm not a real jigglypuff >,< https://www.youtube.com/watch?v=sUKPdN5nSuY

jigglypuff314 (jigglypuff314):

if I have the time, I'll record me playing the song on flute or ocarina :P

OpenStudy (primeralph):

To start, note that: C+F+I >= J and A+D+G <= J and so on. A good way to start is to let C+F+I = X B+E+H = Y A+D+G = Z. So you have X- (Xmod10) = J X-J+Y - (X-J+Y)mod10 = J You see the pattern?

OpenStudy (primeralph):